The function S(t)=139.2-5.1t can be used to estimae the total Cuban sales of windshield wiper fluid, in thousands of dollars, t years after 1990. Predict the total Cuban sales of windshield wiper fluid in 2010.
This problem showed up on a quiz, and still don't know the answer. Please help.

Answer by stanbon(75887)

The function S(t)=139.2-5.1t can be used to estimae the total Cuban sales of windshield wiper fluid, in thousands of dollars, t years after 1990. Predict the total Cuban sales of windshield wiper fluid in 2010.
--------------------
2010 is 20 years after 1990.
S(20) = 139.2 - 5.1(20)
S(20) = 139.2 - 102
S(20) = 37.2
=================
Cheers,
Stan H.